# PRIMAL: Physically Reactive and Interactive Motor Model for Avatar Learning

### v1

@@ 3.4.1 Motion Control with Induced Impulses
A sign of the emergence of physical effects is that the avatar can react to external impulses promptly and naturally. Thus, we can control the avatar with impulses to generate certain actions in a principled manner. For example, we can generate kicks by giving an upward impulse to the joints on the leg, or generate runs by pushing the avatars from the back.
@@ 3.4.2 Motion Control with Classifier-based Guidance
In order to obtain realtime performance, we derive the analytical gradients of these terms instead of using auto-differentiation. Note that we control motions via manipulating $J$ or $\dot{J}$ in the motion state $\bm{x}_{i}$. The SMPL-X parameters, i.e. $(r,\varphi,\theta)$, are not directly controlled.
@@ 4 Model Adaptation
Figure 3: Illustration of our adaptation method. Given a base model, we introduce a ControlNet-like adaptor at each individual transformer block. The control signal $y$ is a generic notation. We adapt the base model for semantic action generation and spatial target reaching. The scale operation is to manipulate the impact of the control signal, corresponding to $\gamma$ in Eq. (3) of CFG.
Since it takes 3-4 days to train our autoregressive diffusion model to obtain high-quality results, we regard it as a base model for task-specific adaptation; this reduces the costs of data and computation. Therefore, we do not need to train from scratch to solve a new task. The effectiveness of this approach has been widely reported in recent advances of foundation models [54, 72, 30, 63, 83, 53].
@@
Few-shot semantic action generation. Given a small set of personalized mocap sequences with action labels, we adapt our base model to generate actions given the labels. In this case, the control signal $y$ shown in Fig. 3 is an integer action label, and an embedding layer maps this label to a latent vector. After adaptation, our model is able to generate infinite personalized movements based on the action labels, and natural transitions between actions, beyond what this small dataset provides. We can control the impact of the action embedding via classifier-free guidance (CFG). Simultaneously, we can also control the movement continuously via the CBG method proposed in Sec. 3.4.2.
Spatial target reaching. In this scenario, the control signal $y$ becomes a 2D target location for the avatar to reach. During finetuning, the pelvis XZ coordinate of the last frame of the canonicalized motion segment is treated as the target. Following [15, 69], we re-scale the target location via $2\cdot(1-e^{-\|\bm{x}\|_{2}})\cdot\frac{\bm{x}}{\|\bm{x}\|_{2}}$, and use a linear layer to map the scaled target location to a latent variable. After training, target-reaching behavior can be synthesized by specifying the target locations at test time. Since the motion is generated autoregressively, one can control the avatar’s motion direction by controlling the target.
@@ 5 Experiments
5.1 Evaluations on Motion Dynamics Learning
A model that learns human motion dynamics well should have two capabilities: 1) It can generate high-fidelity, realistic motions, given an arbitrary initial state; 2) The model’s responses to inputs should be natural and human-like. Thus, we have two sets of experiments to measure motion realism and responsiveness, respectively. Below, we denote our PRIMAL architectures in Fig. 2 as InContext and AdaLN, respectively.
@@ 5.1.1 Motion Realism
Comparisons to state-of-the-art. We compare our methods with two groups of baselines. The off-the-shelf group includes DiP [71], CLoSD [71] and MotionLCM [13]. The “our implemented” group in Tab. 1 includes versions of DiP and DART [89] where the motion representations and pretraining strategies are identical to ours. For details of how we train and fairly compare the methods, see Sup. Mat.
@@ 5.1.2 Motion Responsiveness
Principled Action Generation. Here, we perturb the velocities for subsets of the joints and quantify the resulting motions generated by the agent. Depending on the initial velocities, we can make the agent generate kicks, forward runs, back flips, and forward rolls; see Sup. Mat. for the intuitive principles. With PRIMAL, we only perturb the first frame, but apply the velocities to all history frames of DiP and DART to improve their responsiveness. To measure the motion realism, we generate a 3-second motion for each starting state, leading to 225 sequences per method, and compute ASR, ANCR, and ARD. In addition, we render 30 videos per action for a perceptual study that tests how well the models align with human intuitions; see the Sup. Mat. for details. Table 2 shows that InContext outperforms the prior art. Figure 5 visualizes some results; see Supplementary Video. We see that the avatars have realistic reactions to the velocity interventions.
@@
Table 2: Results of principled action generation.
@@
CBG-based control. The CBG method proposed in Sec. 3.4.2 provides control of the movement direction and speed of the agent. InContext significantly outperforms the baselines on this task – see Sup. Mat. for more details.
5.2 Evaluations on Model Adaptation
For few-shot semantic action generation, we use Mocapade3.0 [19] to capture 5 classes of actions from video (Fig. 6). Although the dataset is only about 3 minutes in total, it is sufficient to fine-tune the model. The quantitative results are shown in Tab. 3. We observe that all these PRIMAL-based methods are effective. The ControlNet approach allows adaptation of the base model to a small set of captured motions, maintaining the controllability and responsiveness of the original model, while being able to mimic the new data.
@@ 6 Conclusion
To create a motor model for embodied intelligence, we propose a new form of autoregressive diffusion model that generates perpetual, controllable, realistic, and responsive motions in real time. Following recent advances in foundation models, our training paradigm has two phases, i.e. 1) unsupervised pretraining to learn motion dynamics, and 2) adaptation to novel tasks. Experimental results show that our model learns motion dynamics, adapts to external velocity inputs, and can be effectively and efficiently adapted to few-shot personalized action generation and spatial target reaching. We demonstrate the advantages with a real-time character animation system built in Unreal Engine.
Limitations and future work. While most motions exhibit physical realism, non-physical motions sometimes occur, e.g. when getting up after falling, the avatar sometimes “flies” up. The method generalizes to falls that it has never seen, but more training data of this type would be helpful. We currently have a limited set of control inputs. Extending this to have more natural inputs through text or speech should be explored. The agent has no awareness of its environment and does not know about obstacles. Learning to condition on the scene is future work. Finally, the model could eventually replace traditional motion graphs and motion matching (cf. [38, 28, 29]) in video games, however, despite running in real time, the computational cost is still too much for this application.
Acknowledgments. We sincerely thank Radu Alexandru Rosu, Yu Sun, and Aman Shenoy for the Rust-based SMPL-X rendering. We are also heartily grateful to Guy Tevet for the help on CLoSD [71]. Moreover, we thank fruitful discussions with Muhammed Kocabas and Nikolas Hesse.
@@ 7.1.2 Test-time Processing
7.1.3 Motion Control with Induced Impulses
@@
kick, left leg left knee, left ankle $(0,0,0.5)$
kick, right leg left knee, left ankle $(0,0,0.5)$
forward run pelvis, neck, shoulders $(0,0,1)$
back flip head $(0,0,-1)$
forward roll head, shoulders, elbows $(0,-0.5,0.5)$
@@ 8.1 Baselines for Motion Realism Evaluation
The “our implemented” group in Tab. 1 contains the DiP and the DART [89] diffusion model. Their motion representations and pretraining strategies are identical to ours. To focus on learning motion dynamics, we remove their text encoders but keep their key designs. Specifically, our DiP version conditions the time embedding via cross attention. Based on 20 frames in the past, it generates 40 future frames. Our DART diffusion model generates 8 future frames based on 2 past frames, suggested in [89]. Neither VAE tokenizers nor scheduled sampling are used. Their training losses are also based on Eq. (4), except that the reconstructed past motions are ignored, following their original settings. During inference, the three test-time processing steps in Sec. 3.4 (or Sec. 7.1.2) are also applied, in order to reduce artifacts such as ground interpenetration and jittering.
8.2 CBG-based Control of Speed and Direction
[…]
@@ 8.3.2 Spatial Target Reaching
We draw 45 initial states from SFU as in Sec. 5.1.2. For each initial state, we specify 8 spatial targets w.r.t. the initial pelvis location, i.e. (±$(\pm{1.5},0),(0,\pm{1.5}),(\pm{1},0),(0,\pm{1})$ in meters. Besides the above two baselines, we run the off-the-shelf DiP model [71] conditioned on the above pelvis goal locations. Each pelvis target also has 45 different motions. We zero mask the text embedding to let the model focus on the spatial target conditioning. We use this DiP model to produce sequences of joint locations, and upsample them to 30 fps via cubic interpolation.
@@ 8.4 Ablation Studies of PRIMAL
Table A5: Ablation studies on the base models. The initial state and every generated motion segments are snapped to the ground. The first part compares different model architectures. The second part shows the influences of the inertialization and joint re-projection. More results are in
8.5 Default Implementations
By default, in PRIMAL the transformer has 10 blocks and the dropout ratio is 0.1. In each transformer block, the self-attention layer has the latent dimension 256 and 8 heads. The feed-forward intermediate layer has 2048 dimensions, and its activation function is SiLU [24, 16, 60]. The DDPM process has 50 steps during training, following [11, 71, 89, 64].
We use AMASS [50] to pretrain our diffusion models  [^3] Our training sets are ACCAD [2], BMLmovi [18], BMLrub [74], CMU [9], DFaust [7], Eyes Japan Dataset [49], HDM05 [51], PosePrior [3], SOMA [17], MoSh [47], SSM.. All sequences are downsampled to 30fps first. During training, 15-frame motion segments are randomly extracted from a batch of sequences at each iteration. We use AdamW [36, 48] optimizer and keep the learning rate at 0.0001. The batch size is fixed to $512$, and the training terminates after 30K epochs, leading to 480K iterations in total. Exponential moving average (EMA) with decay $0.999$ starts to apply after the 1500-th epoch. We use a single NVIDIA H100 GPU for training, taking 3-4 days to pretrain a specific version of PRIMAL.
During inference, we use the checkpoints of EMA. By default, the reverse diffusion process takes 50 steps, and the snapping-to-ground operation (see Sec. 7.1.2) is used.
8.6 Runtime Analysis
@@ 8.7 Perceptual User Studies
We conduct perceptual studies on the Amazon Mechanical Turk platform for both of our experiments mentioned in Sec. 5.1.1 and Sec. 5.1.2. In summary, we present users with paired results—one from our method and one from a baseline method. Users choose the result they prefer, and we report the percentage of cases in which the baseline is preferred.

### v2

@@ 3.4.1 Motion Control with Induced Impulses
The model generalizes outside the training data, exhibiting behavior typically seen only with physics simulation. Specifically, we perturb the avatar with novel external impulses by changing the velocity at one or more joints. Using this, we can control the avatar with impulses to generate certain actions in a principled manner. For example, we can generate kicks by giving an upward impulse to the joints on the leg, or generate runs by pushing the avatars from the back.
@@ 3.4.2 Motion Control with Classifier-based Guidance
In order to obtain real-time performance, we derive the analytical gradients of these terms instead of using auto-differentiation. Note that we control motions via manipulating $J$ or $\dot{J}$ in the motion state $\bm{x}_{i}$. The SMPL-X parameters, i.e. $(r,\varphi,\theta)$, are not directly controlled.
@@ 4 Model Adaptation
Figure 3: Illustration of our adaptation method. Given a base model, we introduce a ControlNet-like adaptor at each individual transformer block. The control signal $y$ is a generic notation. We adapt the base model for semantic action generation and spatial target reaching. The scale operation manipulates the impact of the control signal, corresponding to $\gamma$ in Eq. (3) of CFG.
Since it takes 3-4 days to train our autoregressive diffusion model to obtain high-quality results, we regard it as a base model for task-specific adaptation; this reduces the costs of data and computation. Therefore, we do not need to train from scratch to solve new tasks. The effectiveness of this approach has been widely reported in recent advances of foundation models [53, 72, 30, 63, 83, 52].
@@
Few-shot semantic action generation.
Given a small set of personalized mocap sequences with action labels, we adapt our base model to generate actions given the labels. In this case, the control signal $y$ shown in Fig. 3 is an integer action label, and an embedding layer maps this label to a latent vector. After adaptation, our model is able to generate infinite personalized movements based on the action labels, and natural transitions between actions, beyond what this small dataset provides. We can control the impact of the action embedding via classifier-free guidance (CFG). Simultaneously, we can also perturb the joint velocities to generate natural reactions, and control the movement continuously via the CBG method proposed in Sec. 3.4.2.
Spatial target reaching.
In this scenario, the control signal $y$ becomes a 2D target location for the avatar to reach. During finetuning, the pelvis XZ coordinate of the last frame of the canonicalized motion segment is treated as the target. Following [15, 69], we re-scale the target location via $2\cdot(1-e^{-\|\bm{x}\|_{2}})\cdot\frac{\bm{x}}{\|\bm{x}\|_{2}}$, in order to make the model generalize to targets in extreme scenarios. In addition, we use a linear layer to map the scaled target location to a latent variable. After training, target-reaching behavior can be synthesized by specifying the target locations at test time. Since the motion is generated autoregressively, one can control the avatar’s motion direction by controlling the target.
@@ 5 Experiments
5.1 Evaluations on Body Motion Learning
A model that learns human body motion well should have two capabilities: 1) It can generate high-fidelity, realistic motions, given an arbitrary initial state; 2) The model’s responses to inputs should be natural and human-like. Thus, we have two sets of experiments to measure motion realism and responsiveness, respectively. Below, we denote our PRIMAL architecture in Fig. 2 as InContext. In addition, we compare it with another architecture that exploits adaptive layer normalization [57, 55], denoted as AdaLN, which is illustrated in Sup. Mat.
@@ 5.1.1 Motion Realism
Comparisons to state-of-the-art.
We compare our methods with two groups of baselines. The off-the-shelf group includes DiP [71], CLoSD [71] and MotionLCM [13]. The “our implemented” group in Tab. 1 includes versions of DiP and DART [89] where the motion representations and pretraining strategies are identical to ours. For details of how we train and fairly compare the methods, see Sup. Mat.
@@ 5.1.2 Motion Responsiveness
Table 2: Results of reactions to induced impulses. The evaluation metrics are the same as in Tab. 1. Best results are in boldface.
@@
Reactions to induced impulses.
Here, we perturb the velocities for subsets of the joints and quantify the resulting motions generated by the agent. Depending on the initial velocities, we can make the agent generate kicks, forward runs, back flips, and forward rolls; see Sup. Mat. for the intuitive principles. With PRIMAL, we only perturb the first frame, but apply the velocities to all history frames of DiP and DART to improve their responsiveness. To measure the motion realism, we generate a 3-second motion for each starting state, leading to 225 sequences per method, and compute ASR, ANCR, and ARD. In addition, we render 30 videos per action for a perceptual study that tests how well the models align with human intuitions; see the Sup. Mat. for details. Table 2 shows that InContext outperforms the prior art. Figure 5 visualizes some results; see Supplementary Video. We see that the avatars have realistic reactions to the velocity interventions.
CBG-based control.
The CBG method proposed in Sec. 3.4.2 provides control of the movement direction and speed of the agent. InContext significantly outperforms the baselines on this task – see Sup. Mat. for more details.
@@
5.2 Evaluations on Model Adaptation
For few-shot semantic action generation, we use Mocapade3.0 [19] to capture 5 classes of actions from video (Fig. 6). Although the dataset is only about 3 minutes in total, it is sufficient to fine-tune the model. Quantitative results are shown in Tab. 3, while visualizations are in Fig. 1. We observe that all these PRIMAL-based methods are effective. The ControlNet approach allows adaptation of the base model to a small set of captured motions, maintaining the controllability and responsiveness of the original model, while being able to mimic the new data.
@@ 6 Conclusion
To create a motor model for interactive agents, we propose a new form of autoregressive diffusion model that generates perpetual, controllable, realistic, and responsive motions in real time. Following recent advances in foundation models, our training paradigm has two phases, i.e. 1) unsupervised pretraining to learn body movements, and 2) adaptation to novel tasks. Experimental results show that our model learns motions effectively, adapts to external velocity inputs, and can be efficiently adapted to few-shot personalized action generation and spatial target reaching. We demonstrate the advantages with a real-time character animation system built in Unreal Engine.
Limitations and future work.
First, while most motions exhibit perceptually physical realism, non-physical motions sometimes occur, e.g. when getting up after falling, the avatar sometimes “flies” up. The method generalizes to falls that it has never seen, but more training data of this type would be helpful. We clarify that we do not ensure physical realism. Rather we generate motions that generally look physically realistic. Second, PRIMAL has limited types of control inputs. Extending this to have more natural inputs through text or speech should be explored. For example, we can modify our ControlNet to fuse text and motion features, and fine-tune it on text-motion paired datasets, e.g. BABEL [58] and HumanML3D [20]. Third, the agent has no awareness of its environment and does not know about obstacles. Learning to condition on the scene is future work. Finally, the model could eventually replace traditional motion graphs and motion matching (cf. [38, 28, 29]) in video games, however, despite running in real time (i.e. computation takes less than 0.5 second), the computational cost is still too much for this application.
Acknowledgments.
We sincerely thank Radu Alexandru Rosu, Yu Sun, and Aman Shenoy for the Rust-based SMPL-X rendering. We are also heartily grateful to Guy Tevet for the help on CLoSD [71]. Moreover, we thank fruitful discussions with Muhammed Kocabas and Nikolas Hesse.
@@ Acknowledgments.
7 More Discussions on PRIMAL
7.1 Motivation, Contribution, and Benefits Justification
We aim to build a “motor system” for digital avatars enabling them to move perpetually and react promptly, like real humans. This goes beyond existing game engines that animate characters with canned motions, increasing realism. Inspired by recent advances in motion generation, PRIMAL generates perceptually realistic motions, generalizes across body shapes, supports efficient adaptation, and makes the avatar reactive to impulses. Our experiments and demos present its potential for game production, and show that physical simulation is not necessary to produce character animations that appear physically realistic.
The key novelty is the formulation that generates 0.5-second motion given a single initial state. This contrasts with prior work that generates a long future motion conditioned on a past motion. Its benefits include reducing overfitting, making model training easier, and making the avatar reactive to impulses and classifier-based guidance. Also, our ControlNet’s advantages over existing ones are demonstrated in Sec. 4, experiments, and the SupMat video.
To better understand the benefits of our formulation, we compare two identical settings except the motion length, where ‘ours’ generates 15 frames given 1 frame, and ‘baseline’ generates 40 frames given 20 frames. We replace in-context with cross-attention to handle multi-frame conditioning. Both models are successfully overfit to a ballet sequence with 229 frames, and they can reproduce the ballet conditioned on the start frames. For testing, we first generate 780 future frames given the end frame(s) of that ballet sequence. We find ‘ours’ produces ballet stably, whereas ‘baseline’ gradually fails as time progresses. ASRs (Line475) are 0.08 (‘ours’) and 0.12 (‘baseline’). Second, we generate 156 frames, with conditions from another walking sequence. We find ‘ours’ produces natural transitions to ballet, whereas ‘baseline’ produces severe artifacts. ASRs are 0.06 (‘ours’) vs 0.3 (‘baseline’). These results indicate our setting makes the model more generalizable w.r.t. motion length and semantics. Fig. A1 shows some frames; videos are online on our website.
Figure A1: Our half-second atomic action setting helps generalization and transition.
7.2 Relations to Hierarchical Approaches
Hierarchical approaches like CLoSD [71] and InsActor [62] first use a diffusion-based motion planner to generate a future motion, and then actuate an agent to track the generated motion in simulation. We think this hierarchical setting is not suitable for interactive digital avatars. First, simulation tends to produce less natural looking motion than diffusion-based methods, which is shown by CLoSD’s BPR in Tab. 1 and the Supplementary Video 6:33. Second, simulation can bring extra computational costs and complicate the character animation workflow. PRIMAL only uses the diffusion model, but models much shorter motions. Unlike CLoSD (generates 40 given 20 frames) and InsActor (models more than 100 frames; see their code and SupMat), PRIMAL enables faster avatar responses.
7.3 On Physical Realism Without Simulation
We clarify that we do not ensure physical realism. Rather we generate perceptually realistic motions that look physically realistic. For example, the motion should not contain visible artifacts such as foot skating or ground penetration. Also, the generated motion should be human-like and aligned with user preferences. By learning purely from data, PRIMAL enables physics-like interactions with the agent (e.g. pushing and pulling). It generalizes to disruptions to the first frame that are out of distribution and responds naturally in many situations. As mentioned in the conclusion (Sec. 6), non-physical motions can occur, which could be overcome by using more training data, e.g. of people getting up from the floor.
@@ 8 More Method Demonstrations
8.1 Neural Architecture
We formulate our denoising network as $G(t,\mathbf{X}^{t},\bm{x}_{0})$ with transformers [77, 55], and investigate two architectures that differ in how the initial state is conditioned, as illustrated in Fig. A2. The first architecture leverages in-context conditioning, in which the diffusion time embedding and the initial state embedding are first added and then concatenated with the embedding of the noisy motion segment. We use SiLU [24, 16, 59] in the feed-forward layers. The second architecture follows the DiT network [55] with adaptive layer normalization [57].
Figure A2: The diffusion network is formulated as $\hat{\bm{X}}^{0}=G(t,{\bm{X}}^{t},\bm{x}_{0})$. The tea-green layers contain trainable parameters and the orange blocks are non-learnable operations. The pink squares denote the tokens at individual frames.
@@ 8.2.2 Test-time Processing
8.2.3 Avatar Reactions to Induced Impulses
@@
kick, left leg left knee, left ankle (0,0,0.5)(0,0,0.5)( 0 , 0 , 0.5 )
kick, right leg left knee, left ankle (0,0,0.5)(0,0,0.5)( 0 , 0 , 0.5 )
forward run pelvis, neck, shoulders (0,0,1)(0,0,1)( 0 , 0 , 1 )
back flip head (0,0,1)(0,0,-1)( 0 , 0 , - 1 )
forward roll head, shoulders, elbows (0,0.5,0.5)(0,-0.5,0.5)( 0 , - 0.5 , 0.5 )
@@ 9.1 Baselines for Motion Realism Evaluation
The “our implemented” group in Tab. 1 contains the DiP and the DART [89] diffusion model. Their motion representations and pretraining strategies are identical to ours. To focus on learning the body motion, we remove their text encoders but keep their key designs. Specifically, our DiP version conditions the time embedding via cross attention. Based on 20 frames in the past, it generates 40 future frames. Our DART diffusion model generates 8 future frames based on 2 past frames, suggested in [89]. Neither VAE tokenizers nor scheduled sampling are used. Their training losses are also based on Eq. (4), except that the reconstructed past motions are ignored, following their original settings. During inference, the three test-time processing steps in Sec. 3.4 (or Sec. 8.2.2) are also applied, in order to reduce artifacts such as ground interpenetration and jittering.
@@ 9.3.2 Spatial Target Reaching
[…]
@@
Table A5: Ablation studies on the base models. The initial state and every generated motion segments are snapped to the ground. The first part compares different model architectures. The second part shows the influences of the inertialization and joint re-projection.
9.4 Ablation Studies of PRIMAL
@@
9.5 Default Implementations
By default, in PRIMAL the transformer has 10 blocks and the dropout ratio is 0.1. In each transformer block, the self-attention layer has the latent dimension 256 and 8 heads. The feed-forward intermediate layer has 2048 dimensions, and its activation function is SiLU [24, 16, 59]. The DDPM process has 50 steps during training, following [11, 71, 89, 64]. We use AMASS [50] to pretrain our diffusion models  [^3] Our training sets are ACCAD [2], BMLmovi [18], BMLrub [74], CMU [9], DFaust [7], Eyes Japan Dataset [49], HDM05 [51], PosePrior [3], SOMA [17], MoSh [47], SSM.. All sequences are downsampled to 30fps first. During training, 15-frame motion segments are randomly extracted from a batch of sequences at each iteration. We use AdamW [36, 48] optimizer and keep the learning rate at 0.0001. The batch size is fixed to 512512512, and the training terminates after 30K epochs, leading to 480K iterations in total. Exponential moving average (EMA) with decay 0.9990.9990.999 starts to apply after the 1500-th epoch. We use a single NVIDIA H100 GPU for training, taking 3-4 days to pretrain a specific version of PRIMAL. During inference, we use the checkpoints of EMA. By default, the reverse diffusion process takes 50 steps, and the snapping-to-ground operation (see Sec. 8.2.2) is used.
9.6 Runtime Analysis
@@ 9.7 Perceptual User Studies
We conduct perceptual studies on the Amazon Mechanical Turk platform for both of our experiments mentioned in Sec. 5.1.1 and Sec. 5.1.2. In summary, we present users with paired results—one from our method and one from a baseline method. Users choose the result they prefer, and we report the percentage of cases in which the baseline is preferred. The layouts of the two perceptual studies are shown in Figs. A4 and A5.